OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = (x – 3)2 – 25? Identify the x–intercepts by graphing the quadratic equation.

OpenStudy (anonymous):

\[ y = (x – 3)2 – 25\]\[ y = 2x – 6 – 25\]\[y=2x-31\] can you do it now?

OpenStudy (anonymous):

no im still confused on what that is ... it doesnt give me a graph to grap it on

OpenStudy (anonymous):

You dan't need to, plug in 0 for x to find y intercept and plug 0 for y to find the x intercept.

OpenStudy (cwrw238):

don't you mean (x - 3)^2 - 25?

OpenStudy (anonymous):

maybe, i readf it as w/o a cerokee.

OpenStudy (cwrw238):

the y intercept occurs when x = 0

OpenStudy (cwrw238):

so when x = 0 y = (0-3)^2 - 25 can you work that out?

OpenStudy (anonymous):

i got that one but i havent gotten an of the others yet what is the next step? can you help me

OpenStudy (cwrw238):

do you mean how to get the x-intercepts?

OpenStudy (anonymous):

yes that is what i am talking about

OpenStudy (cwrw238):

ok at the x intercepts y = 0 so you need to solve (x-3)^2 - 25 = 0 adding 25 to both sides (x-3)^2 = 25 can you continue form here?

OpenStudy (anonymous):

thank you

OpenStudy (cwrw238):

yw